A distribution of measurements is relatively mound-shaped with mean 40 and standard deviation 10. 


(a) What approximate proportion of the measurements will fall between 30 and 50? (Enter your answer to two decimal places.) 

(b) What approximate proportion of the measurements will fall between 20 and 60? (Enter your answer to two decimal places.) 

(c) What approximate proportion of the measurements will fall between 20 and 50? (Enter your answer to three decimal places) 

(d) What approximate proportion of the measurements will be greater than 50? (Enter your answer to two decimal places.) 
Answer #1

(c) To find, P( 20 < X < 50 )

Corresponding z values

For X = 20, Z = ( 20 - 40)/10 = -2

For X = 50, Z = (50 - 40)/10 = 1

Thus, P(20 < X < 50) = P(-2 < Z < 1) = 0.819

(d) P( X > 50 )

Corresponding Z = 1

Thus, required proportion = P(Z > 1) = 0.16